A 65 kg student is walking on a slack line, a leng
Answer #1

The weight of the person is balanced by the normal component of the tension
The net forces in the vertical direction is
W - 2 T sin \theta  = 0
Where W is the weight of the person ,T is the tension in the rope , \theta is angle between the tension with the horizontal direction , here \theta = 200
2 T sin 20   = m g
T = mg / 2 sin 20
T = 65 kg x 9.81 m/s2 / 2 x sin 20
T = 932.18 N